# Format as a Prior: Quantifying and Analyzing Bias in LLMs for Heterogeneous Data

### v2

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities across a wide range of natural language tasks (brown2020language). However, their practical deployment remains constrained by key limitations, including factual inaccuracies (commonly referred to as “hallucinations”) (ji2023survey) and incomplete or outdated knowledge (petroni2019language). One promising direction to address these issues is to incorporate external knowledge sources into the reasoning process—allowing models to ground their outputs in more accurate, up-to-date, and contextually relevant information (lewis2020retrieval; gao2023retrieval; huang2022towards).
@@
When homogeneous inputs are converted into heterogeneous ones with equivalent content, accuracy decreases by 9% and 12% on HotpotQA (yang2018hotpotqa) and MuSiQue (trivedi2022musique) (200 samples each), confirming that format heterogeneity can directly impair reasoning performance. This phenomenon may widely arise in heterogeneous reasoning (christmann2024compmix), where key evidence is distributed across texts, tables, infoboxes, and knowledge graphs. LLMs may focus on information from their preferred formats, potentially overlooking crucial data in others. Such bias can result in incomplete or flawed conclusions and undermine the LLMs’ role as impartial and effective synthesizers of heterogeneous inputs.
Although there have been several studies exploring various types of bias in LLMs, such as bias between multi-modal data (zhu2024unraveling; zhang2025evaluating), there is a lack of systematic research on format bias. To address this gap, we present the first comprehensive investigation and analysis of format bias in LLMs. Our study centers on three critical questions: whether such format biases are systematic, which data-level factors contribute to them, and what internal mechanisms in LLMs underlie their emergence.
@@ 2.1 Heterogeneous reasoning
To formalize this challenge, recent benchmarks such as COMPMIX (christmann2024compmix) and CompMix-IR (min2024unihgkr) require cross-format reasoning, stimulating the development of hybrid QA systems that combine structured and unstructured inputs.
Current approaches can be broadly categorized into two types: (1) Unified retrieval frameworks, which abstract away format heterogeneity using shared APIs or embedding spaces (xia2025er; min2024unihgkr); and (2) LLM-centric pipelines, which enhance downstream reasoning via evidence selection, re-ranking, or modular tool use (christmann2024rag; lehmann2024beyond; zhang2024spaghetti; biswal2024text2sql).
@@ 2.2 LLM Behavior under Conflicting Evidence
Recent studies have uncovered a broad range of behavioral biases in LLMs, extending beyond social stereotypes to systematic patterns in reasoning and judgment. A key challenge lies in how LLMs handle conflicts—both between their parametric knowledge (internal beliefs) and in-context evidence, and among competing pieces of evidence (xu2024knowledge). LLMs tend to favor information aligned with their pre-trained knowledge, even when contradicted by accurate inputs (jin2024tug; xie2023adaptive). These LLM biases are shaped by factors such as entity popularity (xie2023adaptive), event recency (fang2023getting), and evidence frequency (jin2024tug). Input artifacts also affect LLM behavior, such as preferring self-generated content over retrieved passages (tan2024blinded). Related efforts have also examined knowledge conflicts in multi-modal (zhang2025evaluating; zhu2024unraveling) and multi-agent settings (ju2025investigating), where preference biases and inter-agent inconsistency further complicate reasoning.
Benchmarks like ConflictBank (su2024conflictbank), WikiContradict (hou2024wikicontradict), and WhoQA (pham2024s) have been proposed to evaluate how LLMs handle factual or semantic inconsistencies, especially in ambiguous scenarios. In response, a range of strategies have emerged, including conflict-aware decoding (yuan2024discerning; jin2024tug), counterfactual data augmentation (fang2023getting), internal intervention via attention pruning (jin2024cutting), neuron reweighting (shi2024ircan), or prompting LLMs to generate multi-answer responses with source attribution (shaier2024adaptive).
@@ 3.1 Dataset and Format Construction
We construct our dataset based on ConflictBank (su2024conflictbank), a public corpus designed to evaluate LLM behavior under factual conflicts. We randomly sample 4,000 entries, each containing one factual claim and three counterclaims with corresponding supporting evidence. This results in 12,000 contradiction pairs, as each claim is paired individually with each of the three counterclaims. Each pair consists of two different claims about the same subject and relation, each supported by its own piece of evidence.
@@ 3.2 Confounding Factor Control
  • Filtering Internal Knowledge: To ensure that LLM responses are based on external evidence rather than parametric memory, we adopt a filtering procedure consistent with prior work (gekhman2024does). Each factual claim is tested 16 times by directly querying a given LLM with the corresponding question in a zero-shot setting, and only those samples for which the model fails to reproduce the factual claim in all trials are retained.
  • Randomizing Evidence Order: To eliminate the known bias introduced by input order (xie2023adaptive), we randomize the sequence of all evidence segments for each input.
@@ 3.3 Evaluated LLMs
This evaluation covers ten LLMs across six major series: GPT-4o-mini (achiam2023gpt), LLaMA-3.1 (8B), Mistral (7B), Qwen3 (8B, 14B, 30B-A3B, 32B) (qwen3technicalreport), Gemma-2 (9B, 27B) (team2024gemma), and GLM-4 (9B) (glm2024chatglmfamilylargelanguage). These LLMs span a range of sizes and architectures, enabling cross-family comparison of format-driven biases.
@@ 3.4 Evaluation Protocol and Metrics
Given the dataset’s scale, we adopt an automated evaluation pipeline using LLMs as adjudicators. Specifically, GPT-4o-mini, GLM-4.5-Air (zeng2025glm), and Qwen-plus are employed to judge which of the two conflicting claims (Source A or Source B) each target response supports.
@@ A.2 Examples of Format Transformation
Following prior work (min2024unihgkr; biswal2024text2sql; christmann2024rag; xia2025er), we directly feed the transformed structured data into the LLMs without additional adapters or format-specific encoders.

### v3

@@ Abstract
Appendixhttps://github.com/NLPGM/Format-as-a-prior/appendix.pdf
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities across a wide range of natural language tasks (Brown et al. 2020). However, their practical deployment remains constrained by key limitations, including factual inaccuracies (commonly referred to as “hallucinations”) (Ji et al. 2023) and incomplete or outdated knowledge (Petroni et al. 2019). One promising direction to address these issues is to incorporate external knowledge sources into the reasoning process—allowing models to ground their outputs in more accurate, up-to-date, and contextually relevant information (Lewis et al. 2020; Gao et al. 2023; Huang and Chang 2022).
@@
When homogeneous inputs are converted into heterogeneous ones with equivalent content, accuracy decreases by 9% and 12% on HotpotQA (Yang et al. 2018) and MuSiQue (Trivedi et al. 2022) (200 samples each), confirming that format heterogeneity can directly impair reasoning performance. This phenomenon may widely arise in heterogeneous reasoning (Christmann et al. 2024), where key evidence is distributed across texts, tables, infoboxes, and knowledge graphs. LLMs may focus on information from their preferred formats, potentially overlooking crucial data in others. Such bias can result in incomplete or flawed conclusions and undermine the LLMs’ role as impartial and effective synthesizers of heterogeneous inputs.
Although there have been several studies exploring various types of bias in LLMs, such as bias between multi-modal data (Zhu et al. 2024; Zhang et al. 2025), there is a lack of systematic research on format bias. To address this gap, we present the first comprehensive investigation and analysis of format bias in LLMs. Our study centers on three critical questions: whether such format biases are systematic, which data-level factors contribute to them, and what internal mechanisms in LLMs underlie their emergence.
@@ 2.1 Heterogeneous reasoning
To formalize this challenge, recent benchmarks such as COMPMIX (Christmann et al. 2024) and CompMix-IR (Min et al. 2024) require cross-format reasoning, stimulating the development of hybrid QA systems that combine structured and unstructured inputs.
Current approaches can be broadly categorized into two types: (1) Unified retrieval frameworks, which abstract away format heterogeneity using shared APIs or embedding spaces (Xia et al. 2025; Min et al. 2024); and (2) LLM-centric pipelines, which enhance downstream reasoning via evidence selection, re-ranking, or modular tool use (Christmann and Weikum 2024; Lehmann et al. 2024; Zhang et al. 2024; Biswal et al. 2024).
@@ 2.2 LLM Behavior under Conflicting Evidence
Recent studies have uncovered a broad range of behavioral biases in LLMs, extending beyond social stereotypes to systematic patterns in reasoning and judgment. A key challenge lies in how LLMs handle conflicts—both between their parametric knowledge (internal beliefs) and in-context evidence, and among competing pieces of evidence (Xu et al. 2024). LLMs tend to favor information aligned with their pre-trained knowledge, even when contradicted by accurate inputs (Jin et al. 2024a; Xie et al. 2023). These LLM biases are shaped by factors such as entity popularity (Xie et al. 2023), event recency (Fang et al. 2023), and evidence frequency (Jin et al. 2024a). Input artifacts also affect LLM behavior, such as preferring self-generated content over retrieved passages (Tan et al. 2024). Related efforts have also examined knowledge conflicts in multi-modal (Zhang et al. 2025; Zhu et al. 2024) and multi-agent settings (Ju et al. 2025), where preference biases and inter-agent inconsistency further complicate reasoning.
Benchmarks like ConflictBank (Su et al. 2024), WikiContradict (Hou et al. 2024), and WhoQA (Pham et al. 2024) have been proposed to evaluate how LLMs handle factual or semantic inconsistencies, especially in ambiguous scenarios. In response, a range of strategies have emerged, including conflict-aware decoding (Yuan et al. 2024; Jin et al. 2024a), counterfactual data augmentation (Fang et al. 2023), internal intervention via attention pruning (Jin et al. 2024b), neuron reweighting (Shi et al. 2024), or prompting LLMs to generate multi-answer responses with source attribution (Shaier et al. 2024).
@@ 3.1 Dataset and Format Construction
We construct our dataset based on ConflictBank (Su et al. 2024), a public corpus designed to evaluate LLM behavior under factual conflicts. We randomly sample 4,000 entries, each containing one factual claim and three counterclaims with corresponding supporting evidence. This results in 12,000 contradiction pairs, as each claim is paired individually with each of the three counterclaims. Each pair consists of two different claims about the same subject and relation, each supported by its own piece of evidence.
@@ 3.2 Confounding Factor Control
  • Filtering Internal Knowledge: To ensure that LLM responses are based on external evidence rather than parametric memory, we adopt a filtering procedure consistent with prior work (Gekhman et al. 2024). Each factual claim is tested 16 times by directly querying a given LLM with the corresponding question in a zero-shot setting, and only those samples for which the model fails to reproduce the factual claim in all trials are retained.
  • Randomizing Evidence Order: To eliminate the known bias introduced by input order (Xie et al. 2023), we randomize the sequence of all evidence segments for each input.
@@ 3.3 Evaluated LLMs
This evaluation covers ten LLMs across six major series: GPT-4o-mini (Achiam et al. 2023), LLaMA-3.1 (8B), Mistral (7B), Qwen3 (8B, 14B, 30B-A3B, 32B) (Team 2025), Gemma-2 (9B, 27B) (Team et al. 2024), and GLM-4 (9B) (GLM et al. 2024). These LLMs span a range of sizes and architectures, enabling cross-family comparison of format-driven biases.
@@ 3.4 Evaluation Protocol and Metrics
Given the dataset’s scale, we adopt an automated evaluation pipeline using LLMs as adjudicators. Specifically, GPT-4o-mini, GLM-4.5-Air (Zeng et al. 2025), and Qwen-plus are employed to judge which of the two conflicting claims (Source A or Source B) each target response supports.
@@ A.2 Examples of Format Transformation
Following prior work (Min et al. 2024; Biswal et al. 2024; Christmann and Weikum 2024; Xia et al. 2025), we directly feed the transformed structured data into the LLMs without additional adapters or format-specific encoders.